OpenStudy (anonymous):

Find S15 for the series -1 + -3 + -5 + -7 +...

OpenStudy (campbell_st):

well you need to common difference... or what is happening to the previous term to get the next. the formula is \[S_{15} = \frac{n}{2}[2a + (n - 1)d]\] you know the number of terms n = 15, the 1st term a = -1 and before you can find the sum, you need the common difference d = substitute the values and evaluate

OpenStudy (anonymous):

i dont understamd

OpenStudy (campbell_st):

Ok... so whats happening in the pattern... -1, -3, -5, ...

OpenStudy (anonymous):

decreasing by 2

OpenStudy (campbell_st):

yep... or -2 so the formula is \[S_{15} = \frac{n}{2}[2a + (n - 1)d]\] and from you information you have a = -1, n = 15 and d = -2 so substituting you get \[S_{15} = \frac{15}{2}\times[2 \times -1 + (15 -1)\times -2]\] just use order of operation to get your answer

OpenStudy (anonymous):

420?

OpenStudy (anonymous):

ANd how would you find common ratio?

OpenStudy (campbell_st):

ummm I think your answer is a little large \[S_{15} = \frac{15}{2}[ -2 -28] \] which is \[S_{15} = \frac{15}{2} [ -30] = 15 \times -15 = -225\]

OpenStudy (campbell_st):

the common ratio occurs in geometric series when terms are multiplied by the same number here is an example 2, 4, 8, 16, 32, .... a term is multiplied by 2 to get the next term an easy way to check is term2/term1 does it equal term3/term2 4/2 = 8/2 both are 2 so the common ratio is 2 hope this helps